# ProRefine: Inference-time Prompt Refinement with Textual Feedback

### v1

@@ Abstract
Agentic workflows, where multiple AI agents collaborate to accomplish complex tasks like reasoning or planning, are becoming increasingly prevalent. However, these workflows often suffer from error propagation and sub-optimal performance, largely due to poorly designed prompts that fail to effectively guide individual agents. This is a critical problem because it limits the reliability and scalability of these powerful systems. We introduce ProRefine, an innovative inference-time prompt optimization method that leverages textual feedback from large language models (LLMs) to address this challenge. ProRefine dynamically refines prompts for multi-step reasoning tasks without additional training or ground truth labels. Evaluated on five benchmark mathematical reasoning datasets, ProRefine significantly surpasses zero-shot Chain-of-Thought baselines by 3 to 37 percentage points. This approach not only boosts accuracy but also allows smaller models to match the performance of larger ones, highlighting its potential for efficient and scalable AI deployment, and democratizing access to high-performing AI.
1 Introduction
A critical condition for the advancement of artificial intelligence (AI) is that AI outputs align with human values [10]. Human feedback, a scarce and subjective resource [11, 14, 26, 44], plays a vital role. LLMs, trained on vast datasets, utilize alignment techniques to generate more human-like and accurate responses [16, 30, 37]. Traditional methods employ Reinforcement Learning from Human Feedback (RLHF) [5], with Proximal Policy Optimization (PPO) [33] being a popular choice for fine-tuning LLMs.
Recent research has explored using LLMs as judges and evaluators [3, 4, 17, 20, 43, 49]. Building on use of LLMs to judge outputs, are compound systems that use LLMs to generate feedback / critiques [27, 31]. Akyurek et al. [1] and Wadhwa et al. [42] explore agentic frameworks to detect, critique, and refine tasks that require factual correctness. Another approach is to iteratively refine output using self-generated feedback [21]. TextGrad [48] introduced automatic “differentiation” via text, using textual feedback to optimize the performance of individual components within a compound AI system.
@@ 1 Introduction
Figure 1: Overview of the ProRefine system, illustrating the iterative process of prompt optimization using feedback from LLMs in an agentic workflow.
@@
Our work focuses on optimizing the prompt, a key element in chain-of-thought (CoT) [45] based LLM reasoning. CoT mimics human problem-solving by breaking down complex tasks into smaller, manageable steps. The system we introduce here, ProRefine (Inference-time Prompt Refinement with Textual Feedback), builds upon CoT by adaptively improving prompts using feedback ($LLM_{feedback}$) and an optimizer ($LLM_{optimizer}$) to refine prompts for the task-performing LLM ($LLM_{task}$). This workflow (Figure 1), motivated by the teacher-student framework [39] where a teacher agent guides a student agent to perform a task by providing feedback at intermediate steps, but implemented via LLM interactions without pre-training, represents a novel approach to adaptive agentic reasoning. We explore policy optimization for aligning compound AI systems, drawing inspiration from TextGrad and policy gradient algorithms like PPO. This work investigates the following research questions.
@@
Can textual feedback be used to improve the performance of LLMs at inference-time?
@@
Does model size impact the ability of LLMs to utilize textual feedback?
@@
What is the impact of using a verifier at inference-time?
While prior work has explored prompt optimization [7, 8, 34, 46], they all require training examples to produce task-specific prompts. By contrast, ProRefine is task agnostic and requires no additional training or ground-truth labels. ProRefine is an inference-time optimization method that relies on test-time compute availability and the ability of LLMs to provide and act upon feedback for optimization, see Figure 2 for an example. This offers a valuable alternative to solely relying on scaling up the base model size for all queries, especially in scenarios where the smaller model suffices for most tasks but needs enhancement for specific critical ones.
ProRefine’s ability to break complex tasks into smaller steps and self-improve using feedback makes it a practical solution for multi-step agentic workflows and resource-constrained environments where deploying the largest models isn’t feasible, but temporary access to a capable feedback LLM (perhaps via a separate API call) is possible for critical tasks. This method is also suitable for black-box LLMs, where the internal parameters are inaccessible and only API access is available. ProRefine could prove to be crucial in situations demanding higher interpretability, where the feedback steps (outputs of $LLM_{feedback}$) offer insights into the reasoning correction process, and applications requiring dynamic adaptation without retraining/fine-tuning cycles. To demonstrate the effectiveness of ProRefine, we evaluate our method across five benchmark mathematical reasoning datasets.
@@
  • We propose a novel method - ProRefine for inference-time prompt optimization using textual feedback.
@@
  • We evaluate ProRefine on five datasets: object counting, word sorting, grade-school math problem-solving, math word problems, and algebraic word problems, and compare our method against CoT and TextGrad.
  • We evaluate the importance of using a verifier at inference-time.
Figure 2: ProRefine example. Given an input query (which in this case has correct answer: 13), and an initial prompt, the task model ($LLM_{task}$) gives an incorrect answer. Prorefine uses two additional models, $LLM_{feedback}$ and $LLM_{optimizer}$, to iteratively improve the prompt as the $LLM_{task}$ generates its response. Refining the prompt during generation allows the feedback model to target local regions of the response, providing finer-grained feedback than waiting for the response to complete.
2 Related Work
@@
ProRefine draws inspiration from and contributes to several interconnected research areas. The performance of LLMs is heavily dependent on the quality of the prompts they receive. Early efforts in this domain centered on manual prompt crafting [45], a meticulous process of designing effective prompts to elicit desired responses. Recognizing the limitations and scalability challenges of manual methods, research has increasingly focused on automatic prompt optimization.
@@
Some pioneering automatic methods, such as AutoPrompt [34] and RLPrompt [7] , employ gradient-based search and reinforcement learning techniques, respectively. AutoPrompt [34] for instance, uses gradient-based search to generate prompts for masked language models. It reformulates tasks as fill-in-the-blank problems, achieving performance comparable to supervised models in tasks like sentiment analysis. However, it requires training data and gradient access, limiting its applicability to black-box models. Other approaches leverage LLMs themselves for prompt generation [22, 27, 46, 47, 50]. ProRefine distinguishes itself by operating at inference-time without training data or ground truth labels, extending the concept of Chain-of-Thought prompting [45] by iteratively refining the prompt.
There is a substantial and growing body of work exploring the capacity of LLMs to act as judges or evaluators [3, 4, 17, 20, 41, 43, 49, 51] of text quality, correctness, and other attributes. This capability is crucial for automating various aspects of model development and deployment, from assessing generation quality to providing feedback for model improvement. ProRefine utilizes this concept, employing an LLM to provide feedback (a form of evaluation), but crucially uses this feedback for optimization and combines in a novel fashion inference-time operations with LLM-generated feedback.
@@
The idea of LLM self-improvement and iterative refinement is also highly relevant. Self-Refine [21] is a prominent example, where an LLM generates both output and feedback, using the latter for refinement. Other works explore self-critiquing [31] and reinforcement learning for critique generation (RL4F) [1], along with various feedback and refinement mechanisms [8, 15, 28, 29, 32, 42]. ProRefine shares this spirit but focuses specifically on prompt refinement, making it suitable for agentic workflows and black-box LLMs, unlike direct output modification.
ProRefine builds upon the core ideas in TextGrad [48], which uses textual feedback for optimization within compound AI systems. However, ProRefine applies feedback at intermediate steps for inference-time prompt optimization and rapidly evolving tasks where fine-tuning is infeasible, a key distinction. Moreover, TextGrad is a supervised fine-tuning method, while ProRefine requires no training data, making its data requirements less onerous. Finally, ProRefine belongs to the broader category of inference-time methods [24, 35], that improve LLMs without weight modification, such as multi-agent debate [9]. Inference-time methods aim to improve the performance of models by utilizing test-time compute resources. Recent advances in understanding reasoning in LLMs [12, 13] also provide context to our prompt improvement.
In essence, ProRefine uniquely integrates inference-time operation, LLM-generated feedback, and prompt optimization to enhance multi-step reasoning in agentic workflows, effectively addressing limitations of prior work related to reliance on training data and applicability to black-box models.
3 ProRefine
@@
$LLM_{feedback}$: A model that critiques the $LLM_{task}$’s output, providing detailed feedback on improvements. This model should be capable of providing insightful and accurate critiques [2, 31].
@@
Input: Query: $q$, Initial prompt: $p$, tokens_per_step: $k$, max_steps: $n$, LLMs: $LLM_{task}$, $LLM_{feedback}$, $LLM_{optimizer}$
@@
       $o_{i}=LLM_{task}(p^{*},q)$ // Generate $i*k$ tokens
@@
  • Generation: Use $LLM_{task}$ to generate an output based on the current prompt $p^{*}$ and query $q$. This step is limited to $k$ tokens to control the granularity of feedback.
@@
3.1 Unifying Verifier and Feedback
At inference time, verifiers play a crucial role in judging model outputs [6, 18, 36]. Often trained on extensive corpora, these verifiers can be outcome-based, assessing the final result’s correctness, or process-based, evaluating the validity of intermediate steps [18, 36, 40]. For simplicity in this study, we do not train a bespoke verifier; rather, we employ the Llama3.1-70B-instruct model to function as both the feedback mechanism ($LLM_{feedback}$) and the verifier. A smaller model, specifically fine-tuned for these tasks, could also be used. The verifier’s function is to evaluate the initial output generated by $LLM_{task}$ for each query. If this output is assessed as incorrect, the refinement process is triggered; otherwise, the output is used directly.
@@ 3.2 Implementation Details
4 Experiments and Evaluation
4.1 Data
We evaluate ProRefine on five reasoning tasks, each of which involves multi-step reasoning, making them suitable for evaluating prompt optimization in agentic workflows: object counting and word sorting from the BIG-Bench Hard benchmark [38], grade-school math problem-solving from GSM8K [6], math word problems from SVAMP [25], and algebraic word problems from AQUARAT [19]. We use the same dataset splits and evaluation as Yuksekgonul et al. [48] for object counting, word sorting, and GSM8K. Refer to Table 1 for dataset statistics.
Dataset Train Validation Test
Object Counting 50 100 100
Word Sorting 50 100 100
GSM8K 200 300 1319
SVAMP 2516 622 1000
AQUARAT 97467 254 254
Table 1: Datasets used for evaluating ProRefine on multi-step reasoning tasks.
@@ 4.1 Data
4.2 Experimental Setup
We experiment with three models - Llama3.2-1B-instruct, Llama3.2-3B-instruct, and Llama3.1-8B-instruct [23] for $LLM_{task}$. The prompts are optimized using Algorithm 1, with Llama3.1-70B-instruct used for feedback generation, prompt optimization, and evaluation. We select the values of hyperparameters $k=10$ and $n=25$ to control the granularity of feedback and duration of optimization. Hyperparameters $k$ and $n$ were fixed based on general preliminary exploration and not tuned per task using benchmark training/validation data.
We compare the performance of our method against the zero-shot Chain-of-Thought (CoT) baseline and TextGrad [48], and report test accuracy with 95% confidence interval. We choose TextGrad as a baseline because Yuksekgonul et al. [48] reported performance at par or better than DSPy [15] for prompt optimization on object counting, word sorting, and GSM8k datasets. It is essential to remember that TextGrad is a supervised fine-tuning method that utilizes both the training and validation sets. For TextGrad, we use a comparative setup consisting of a task model to be fine-tuned and Llama3.1-70B-instruct model for feedback generation and backpropagation. The results are shown in Table 2 and Figure 3.
4.3 Results

### v2

@@ Abstract
Agentic workflows, where multiple AI agents collaborate to accomplish complex tasks like reasoning or planning, play a substantial role in many cutting-edge commercial applications, and continue to fascinate researchers across nearly all fields for their potential to accomplish expensive, complex tasks that, until recently, only humans have been trusted to do. These workflows depend critically on the prompts used to provide the roles models play in such workflows. Poorly designed prompts that fail even slightly to guide individual agents can lead to sub-optimal performance that may snowball within a system of agents, limiting their reliability and scalability. To address this important problem of inference-time prompt optimization, we introduce ProRefine, an innovative inference-time optimization method that uses an agentic loop of LLMs to generate and apply textual feedback. ProRefine dynamically refines prompts for multi-step reasoning tasks without additional training or ground truth labels. Evaluated on five benchmark mathematical reasoning datasets, ProRefine significantly surpasses zero-shot Chain-of-Thought baselines by 3 to 37 percentage points. This approach not only boosts accuracy but also allows smaller models to approach the performance of their larger counterparts. This highlights its potential for building more cost-effective and powerful hybrid AI systems, thereby democratizing access to high-performing AI.
Introduction
The rapid evolution of Large Language Models (LLMs) has transformed applications from search engines to automated reasoning systems (Feng et al. 2024). Yet, their efficacy criticality depends on alignment with human values and preferences, a process often bottlenecked by the scarcity and subjectivity of human feedback (Gray and Suri 2019; Kahneman, Sibony, and Sunstein 2021; Prabhakar, Griffiths, and McCoy 2024; Weerasooriya et al. 2023). LLMs, trained on vast datasets, utilize alignment techniques to generate more human-like and accurate responses (Kirk et al. 2024; Rao et al. 2023; Sorensen et al. 2024).
@@ Introduction
Figure 1: Overview of ProRefine system, illustrating the iterative process of prompt optimization using feedback from LLMs. In each iteration, $LLM_{task}$ extends its output by an additional k tokens, enabling step-by-step feedback to progressively refine the prompt with $LLM_{optimizer}$.
Recent advances have shifted toward leveraging LLMs themselves as judges and evaluators (Bavaresco et al. 2024; Chiang and Lee 2023; Li et al. 2024; Liu et al. 2023; Wang et al. 2023; Zheng et al. 2023). Building on this, compound AI systems that use LLMs to generate feedback / critiques (Pryzant et al. 2023; Saunders et al. 2022). Akyurek et al. (2023) and Wadhwa et al. (2024) explore agentic frameworks to detect, critique, and refine tasks that require factual accuracy. Another approach is to iteratively refine the output using self-generated feedback (Madaan et al. 2023). TextGrad (Yuksekgonul et al. 2024) introduced automatic “differentiation” via text, using textual feedback to optimize the performance of individual components within a compound AI system.
@@
Our work focuses on optimizing the prompt, a key element in chain-of-thought (CoT) (Wei et al. 2022) based LLM reasoning. CoT mimics human problem-solving by breaking down complex tasks into smaller, manageable steps. The system we introduce here, ProRefine (Inference-time Prompt Refinement with Textual Feedback), builds upon CoT by adaptively improving prompts using feedback ($LLM_{feedback}$) and an optimizer ($LLM_{optimizer}$) to refine prompts for the task-performing LLM ($LLM_{task}$). This workflow (Figure 1), motivated by the teacher-student framework (Torrey and Taylor 2013) where a teacher agent guides a student agent to perform a task by providing feedback at intermediate steps, but implemented via LLM interactions without pre-training, represents a novel approach to adaptive agentic reasoning. We explore policy optimization for aligning compound AI systems, drawing inspiration from TextGrad and policy gradient algorithms like PPO. This work investigates the following research questions.
@@
How effectively can textual feedback enhance LLM performance of LLMs during inference?
@@
To what extent does model size impact the ability of LLMs to utilize textual feedback?
@@
What is the impact of incorporating a verifier on accuracy at inference time?
Figure 2: ProRefine example. Given an input query (which in this case has correct answer: 13), and an initial prompt, the task model ($LLM_{task}$) gives an incorrect answer. ProRefine uses two additional models, $LLM_{feedback}$ and $LLM_{optimizer}$, to iteratively improve the prompt as the $LLM_{task}$ generates its response. Refining the prompt during generation allows the feedback model to target local regions of the response, providing finer-grained feedback than waiting for the response to complete. We also provide an additional example illustrating our approach in Figure 5.
Although prior work has explored prompt optimization (Deng et al. 2022; Dong et al. 2024; Shin et al. 2020; Yang et al. 2024), they all often focus on either offline fine-tuning, which requires extensive training data, or universal application of largest, most capable models to every task. This presents a practical dilemma in many real-world scenarios. Continuously fine-tuning is not always feasible, and relying exclusively on state-of-the-art models is often computationally prohibitive. A different approach is needed for scenarios that require dynamic, on-the-fly repair for specific and difficult queries where a standard prompt fails. This is particularly true in resource-aware deployments, where a smaller model may suffice for most tasks but requires enhancement for a small subset of critical queries. The goal, therefore, shifts from finding a single, universally optimal prompt to performing targeted, inference-time intervention.
This hybrid-model paradigm makes a method like ProRefine a practical solution. It is designed for resource-constrained environments where deploying the largest models for every query isn’t feasible, but temporary access to a capable feedback LLM (perhaps via a separate API call) is possible for critical tasks. In such cases, the refinement process is triggered as an on-demand “expert intervention.” ProRefine is task-agnostic and requires no additional training or ground-truth labels. It is an inference-time optimization method that relies on the availability of test-time compute and the ability of LLMs to provide and act upon feedback for optimization.
The ability to break complex tasks into smaller steps and dynamically improve prompts offers a crucial advantage in multi-step agentic workflows where errors can compound. As illustrated in Figure 2, this method is also suitable for black-box LLMs where only API access is available. ProRefine could prove to be crucial in situations demanding greater interpretability, where feedback steps (outputs of $LLM_{feedback}$) offer insights into the reasoning correction process and applications requiring dynamic adaptation without retraining/fine-tuning cycles. To demonstrate its effectiveness, we evaluate ProRefine across five benchmark mathematical reasoning datasets, showing it offers a robust alternative to solely scaling up the base model for all queries.
@@
  • We propose a novel method - ProRefine - for prompt optimization inference-time using textual feedback.
@@
  • We evaluated ProRefine on five datasets: object counting, word sorting, grade-school math problem solving, math word problems, and algebraic word problems, and compared our method against CoT and TextGrad.
  • We evaluate the importance of using a verifier at inference time.
Essentially, ProRefine is an elegant, training-free, inference-time solution that fixes the core prompt to improve reasoning, making it suitable for practical hybrid-model deployment.
Related Work
@@
ProRefine draws inspiration from and contributes to several interconnected research areas. The performance of LLMs is heavily dependent on the quality of the prompts they receive. Early efforts in this domain centered on manual prompt crafting (Wei et al. 2022), a meticulous process of designing effective prompts to elicit desired responses. Recognizing the limitations and scalability challenges of manual methods, research has increasingly focused on automatic prompt optimization with a growing emphasis on agentic workflows that enable dynamic and adaptive reasoning.
Prompt Generation
@@
Some pioneering automatic methods, such as AutoPrompt (Shin et al. 2020) and RLPrompt (Deng et al. 2022), employ gradient-based search and reinforcement learning techniques, respectively. AutoPrompt (Shin et al. 2020) uses gradient-based search to generate prompts for masked language models. It reformulates tasks as fill-in-the-blank problems, achieving performance comparable to supervised models in tasks like sentiment analysis. However, it requires training data and gradient access, limiting its applicability to black-box models. Other approaches leverage LLMs themselves for prompt generation (Mehta, Pyatkin, and Srikumar 2024; Pryzant et al. 2023; Yang et al. 2024, 2022; Zhou et al. 2022). Recent works like Promptomatix (Murthy et al. 2025) and EvoAgentX (Wang et al. 2025) extend this direction by enabling automatic prompt refinement across multiple tasks, workflows, and tools. ProRefine distinguishes itself by operating simply at inference-time, requiring no training data, gradient access, or model retraining, while enabling prompt refinement in dynamically evolving settings.
Self-Refinement
There is a substantial and growing body of work exploring the capacity of LLMs to act as judges or evaluators (Bavaresco et al. 2024; Chiang and Lee 2023; Li et al. 2024; Liu et al. 2023; Verga et al. 2024; Wang et al. 2023; Zheng et al. 2023; Zhuge et al. 2024). This capability has been leveraged to assess response quality or provide self-feedback. ProRefine adopts this principle, using LLM-generated textual feedback to improve its own prompting process. Unlike prior uses of LLM evaluation solely for ranking or filtering, ProRefine uses that feedback in a closed-loop for optimization during inference.
@@
The idea of LLM iterative refinement is highly relevant. Self-Refine (Madaan et al. 2023) is a prominent example, where an LLM generates both output and feedback, using the latter for refinement. ARIES (Zeng et al. 2025) further enhances refinement via Elo-style agent debate. Other works explore self-critiquing (Saunders et al. 2022) and reinforcement learning for critique generation (RL4F) (Akyurek et al. 2023), along with various feedback and refinement mechanisms (Dong et al. 2024; Khattab et al. 2024; Qu et al. 2024; Ranaldi and Freitas 2024; Schick et al. 2023; Wadhwa et al. 2024), and Monte Carlo-based refinement in math reasoning (MC-NEST) (Rabby, Keya, and Auer 2025). While ProRefine shares the self-refinement spirit, it focuses on prompt refinement, suitable for agentic workflows and black-box LLMs, while avoiding reinforcement learning and direct output modification.
Inference-Time Scaling
ProRefine belongs to the broader category of inference-time methods (Muennighoff et al. 2025; Snell et al. 2024a), that improve LLMs without without weight modification (Du et al. 2024). Inference-time methods aim to improve the performance of models by utilizing test-time compute resources. TextGrad (Yuksekgonul et al. 2024) performs gradient-free inference-time optimization using textual feedback. ProRefine applies a similar idea to intermediate prompt refinement for dynamic reasoning chains. TextGrad relies on supervised fine-tuning, whereas ProRefine operates without training data, offering ease of integration. Other inference-time strategies include RL-of-Thoughts (Hao et al. 2025) and Reward-Is-Enough (Song et al. 2025), which apply RL-based signal propagation during inference. AvR (Alignment via Refinement) (Zhang et al. 2025a) proposes recursive CoT refinement using long-form reasoning. ProRefine, by contrast, performs step-level feedback on prompts rather than final outputs, and requires no external tools or supervision.
Agentic Workflows
ProRefine also fits into a broader trend toward agentic workflows. AFlow (Zhang et al. 2025b) automates agentic workflows through prompt-based search over prior structures, while EvoAgentX (Wang et al. 2025) evolves agent behaviors and topologies. Meanwhile, Mass (Zhou et al. 2025) and DebFlow (Su et al. 2025) optimize multi-agent configurations via interleaved search and debate. ProRefine focuses instead on optimizing individual agent prompts within fixed workflows, complementing these methods. Unlike tool-integrated or debate-based systems, ProRefine remains model-agnostic and easy to integrate into any prompt-based agent loop.
ProRefine
@@
$LLM_{feedback}$: A model that critiques the $LLM_{task}$’s output, providing detailed feedback on improvements. This model should be capable of providing insightful and accurate critiques (Bai et al. 2022; Saunders et al. 2022).
@@
Input: Query: $q$, Initial prompt: $p$, tokens_per_step: $k$, max_steps: $n$, LLMs: $LM_{task}, LL$$M_{feedback}, LL$$LM_{optimizer}$
@@
 $o_{i}=LLM_{task}(p^{*},q)$ // Generate iki*k tokens
@@
  • Generation: Use $LLM_{task}$ to generate an output based on the current prompt $p^{*}$ and query $q$. This step is limited to $i$ *$k$ tokens to control the granularity of the feedback. In each iteration, $LLM_{task}$ produces $k$ more tokens, attempting to refine prior output while progressively continuing its response to the query.
@@
Unifying Verifier and Feedback
At inference time, verifiers play a crucial role in judging model outputs (Cobbe et al. 2021; Lightman et al. 2024; Snell et al. 2024b). Often trained on extensive corpora, these verifiers can be outcome-based, assessing the final result’s correctness, or process-based, evaluating the validity of intermediate steps (Lightman et al. 2024; Snell et al. 2024b; Uesato et al. 2022). For simplicity in this study, we do not train a bespoke verifier; rather, we employ the Llama3.1-70B-instruct model to function as both the feedback mechanism ($LLM_{feedback}$) and the verifier. We manage these roles through separate API calls, each with a role-defining prompt. A smaller model, specifically fine-tuned for these tasks, could also be used. The verifier’s function is to evaluate the initial output generated by $LLM_{task}$ for each query. If verifier assess this output is incorrect, the refinement process is triggered; otherwise, the output is used as is. This also saves computation on answers that are already correct.
@@
Experiments and Evaluation
Data
We evaluate ProRefine on five reasoning tasks, each of which involves multi-step reasoning, making them suitable for evaluating prompt optimization in agentic workflows. We include the original dataset split sizes in (train/validation/test) format: object counting and word sorting from the BIG-Bench Hard benchmark (Srivastava et al. 2023) (50/100/100), grade-school math problem-solving from GSM8K (Cobbe et al. 2021) (200/300/1319), math word problems from SVAMP (Patel, Bhattamishra, and Goyal 2021) (2516/622/1000), and algebraic word problems from AQUARAT (Ling et al. 2017) (97467/254/254). We use the same splits and evaluation as Yuksekgonul et al. (2024) for object counting, word sorting, and GSM8K.
Experimental Setup
@@
Figure 4: Average number of prompt refinement iterations.
We experiment with three models - Llama3.2-1B-instruct, Llama3.2-3B-instruct, and Llama3.1-8B-instruct (Meta 2024) for $LLM_{task}$. The prompts are optimized using Algorithm 1, with Llama3.1-70B-instruct used for feedback generation, prompt optimization, and evaluation. We select the values of hyperparameters $k=10$ and $n=25$ to control the granularity of feedback and duration of optimization. Hyperparameters $k$ and $n$ were fixed based on general preliminary exploration and not tuned per task using benchmark training/validation data.
We compare the performance of our method against the zero-shot Chain-of-Thought (CoT) baseline and TextGrad (Yuksekgonul et al. 2024), and report test accuracy with 95% confidence interval. We choose TextGrad as a baseline because Yuksekgonul et al. (2024) reported performance at par or better than DSPy (Khattab et al. 2024) for prompt optimization on object counting, word sorting, and GSM8k datasets. It is essential to remember that TextGrad is a supervised fine-tuning method that utilizes both the training and validation sets. For TextGrad, we use a comparative setup consisting of a task model to be fine-tuned and Llama3.1-70B-instruct model for feedback generation and backpropagation. The results are shown in Table 1 and Figure 3.
Results